### 06 Mar 2026 v1

@@ 5 Numerical Results
In this section, we present numerical experiments to assess the performance of the proposed LAE-EnKF in nonlinear data assimilation problems. The experiments are designed to evaluate three key aspects of the method: assimilation accuracy, robustness over time, and the benefit of enforcing linear dynamics in a learned latent space. All of the following numerical examples are carried out on a computing server with a NVIDIA GeForce RTX 3090 GPU.
To enrich the information content of partial observations, we adopt a time-delay embedding strategy [32]. Specifically, a sequence of $L$ consecutive observations is concatenated to form an augmented observation vector,
@@
which is then processed by the observation encoder $\mathscr{E}_{\mathrm{obs}}$. This approach allows the latent observation embedding to exploit short-term temporal correlations without modifying the filtering formulation.
@@
(15) $\mathrm{E}_{1:T}=\sqrt{\frac{1}{TD}\sum_{k=1}^{T}\sum_{i=1}^{D}\big(\widehat{x}_{k,i}-x_{k,i}\big)^{2}},\qquad\mathrm{E}_{\mathrm{Rel},1:T}=\frac{\mathrm{E}_{1:T}}{\sqrt{\frac{1}{T}\sum_{k=1}^{T}\left(\frac{1}{D}\sum_{i=1}^{D}x_{k,i}^{2}\right)}}.$
@@
    EnKF: the standard ensemble Kalman filter, where both forecast and analysis steps are carried out directly in the physical state space. Compared with other methods, the implementation of EnKF would require the usage of the forecast and observation models.
@@
    AE-EnKF: an autoencoder-assisted EnKF that uses a nonlinear autoencoder for dimensionality reduction, but performs the Kalman update in the physical space and does not impose any structure on the latent dynamics.
@@
    DAE-EnKF: a deep autoencoder-based latent assimilation method, in which the EnKF update is performed in a learned latent space and the latent dynamics are modeled by a general neural network without linear constraints. This setting is similar in spirit to [24]. To ensure a fair comparison, the encoder and decoder architectures are identical to those used in the proposed LAE framework.
All methods use the same ensemble size, data normalization, and training configuration to ensure comparability. In particular, all state variables and observations are normalized to zero mean and unit variance prior to training and assimilation. In all experiments, the encoder and decoder are implemented using convolutional neural networks(CNN), which are well suited for high-dimensional systems with spatial structure. The state encoder consists of three convolutional layers. Each layer uses a $4\times 4$ kernel with stride $2$, progressively reducing the spatial resolution while extracting multiscale features. A fully connected layer with LeakyReLU activations (negative slope $\alpha=0.1$) then maps the resulting features to the latent variables. The decoder mirrors this architecture and reconstructs the physical state using transposed convolutional layers. The latent linear transition operator $\mathbf{A}$ is learned as a single fully connected neural network without any activation function. This architectural choice explicitly enforces linear latent dynamics and is essential for maintaining structural compatibility with Kalman filtering. The observation encoder is implemented as a lightweight network composed of two one-dimensional convolutional layers with kernel size $3$, followed by a fully connected layer that projects the observations into the latent space. This design allows observations with different spatial layouts to be embedded consistently with the latent state representation. The loss weights in (3.2) are selected so that the reconstruction, prediction, latent consistency, and regularization terms contribute at comparable magnitudes during training. All networks are implemented in PyTorch and trained using the Adam optimizer with an initial learning rate of $10^{-3}$. Early stopping based on the validation RMSE is applied to mitigate overfitting.
@@ 5.1 Toy example
The system is governed by a single angular variable $\theta_{t}$, while the physical state is defined as
(16) $\mathbf{x}_{t}=W\begin{bmatrix}\cos(\theta_{t})\\ \sin(\theta_{t})\end{bmatrix}\in\mathbb{R}^{100},\qquad W\in\mathbb{R}^{100\times 2},\quad W_{ij}\sim\mathcal{N}(0,1),$
which embeds a two-dimensional rotational motion into a 100-dimensional ambient space. The angular dynamics evolve according to
(17) $\theta_{t+1}=\theta_{t}+\delta+\alpha\sin(2\theta_{t})+c\,\epsilon_{t},$
where $\delta=\pi/50$ is a constant rotation increment, $\alpha=0.01$ introduces weak nonlinearity, $c=0.01$ scales the process noise, and $\epsilon_{t}\sim\mathcal{N}(0,1)$.
@@
(18) $\mathbf{y}_{t}=H\mathbf{x}_{t}+\boldsymbol{\eta}_{t},\qquad\boldsymbol{\eta}_{t}\sim\mathcal{N}(0,0.1^{2}),$
where $H\in\mathbb{R}^{2\times 100}$ randomly selects two state components. We simulate $N_{\mathrm{traj}}=500$ independent trajectories of length $T=100$, with initial angles $\theta_{0}$ sampled uniformly from $\pi,\pi]$. The data are split into $90\%$ for training and $10\%$ for testing.
@@
Figure 4: Long-term prediction relative error for latent dimensions $n=2,3,4$. Predictions are obtained by recursively propagating the learned latent dynamics forward in time and decoding back to the physical space.
Finally, we evaluate data assimilation performance using LAE-EnKF and compare it with EnKF, AE-EnKF, and DAE-EnKF. Tab. 1 reports the global relative RMSE $\mathrm{E}_{\mathrm{Rel},1:T}$ for latent dimensions $n=2,3,4$. Fig. 5 shows the time evolution of the reconstruction error for $n=2$, together with $95\%$ confidence intervals over ten independent runs. While all learned-model-based methods outperform the standard EnKF, AE-EnKF and DAE-EnKF exhibit noticeable temporal fluctuations. In contrast, LAE-EnKF produces smoother and more stable error trajectories.
[…]
Figure 5: Time evolution of the relative RMSE for different methods with latent dimension $n=2$. Solid curves show the mean over 10 independent runs, and shaded regions indicate the corresponding 95% confidence intervals.
[…]
Table 1: Relative error $\mathrm{E}_{\text{Rel},1:T}$ for different methods with various values of $n$.
@@ 5.2 Advection–diffusion–reaction equation
We consider a nonlinear advection diffusion reaction equation on the two-dimensional spatial domain $\mathbf{x}=(x_{1},x_{2})\in[0,1]\times[0,1]$,
(19) $\partial_{t}u(\mathbf{x},t)+\mathbf{v}(\mathbf{x})\cdot\nabla u(\mathbf{x},t)$ $=\mu\Delta u(\mathbf{x},t)-\alpha\bigl(\sin u(\mathbf{x},t)-u(\mathbf{x},t)\bigr),$
$u(\mathbf{x},0)$ $=u_{0}(\mathbf{x}),$
subject to periodic boundary conditions. The diffusion coefficient and reaction parameter are set to $\mu=10^{-3}$ and $\alpha=0.8$, respectively. The advection velocity field is given by
$\mathbf{v}(\mathbf{x})=\bigl(\sin(2\pi x_{2}),\,\sin(2\pi x_{1})\bigr).$
@@
(20) $u(\mathbf{x},0)\sim\mathcal{GP}\bigl(0,K_{\ell}(\mathbf{x},\mathbf{x}^{\prime})\bigr),$
@@
$K_{\ell}(\mathbf{x},\mathbf{x}^{\prime})=\exp\!\left(-\frac{\sin^{2}(\pi(x_{1}-x_{1}^{\prime}))+\sin^{2}(\pi(x_{2}-x_{2}^{\prime}))}{2\ell^{2}}\right).$
Data assimilation is performed using sparse and noisy observations of the true solution. Measurements are collected at $5\times 5$ uniformly distributed sensor locations, resulting in an observation vector $\mathbf{y}_{k}\in\mathbb{R}^{25}$. Observations are available every $\Delta t=0.05$ up to the final time $T=1.5$. Fig. 7 shows the true solution at the initial and final times, with sensor locations marked by white circles. The observation model is
(21) $\mathbf{y}_{k}=\mathbf{y}^{\mathrm{true}}_{k}+\boldsymbol{\eta}_{k},\qquad\boldsymbol{\eta}_{k}\sim\mathcal{N}\!\left(\mathbf{0},\,\sigma^{2}\mathbf{I}\right),$
where the noise level is set to $\sigma=0.01$. The initial ensemble is generated by sampling spatial white noise followed by convolution with a two-dimensional Gaussian filter with standard deviation $3$. This procedure introduces spatial correlation and provides a common prior ensemble for all filtering methods.
@@
Figure 7: True solution fields of the advection–diffusion–reaction equation at the initial time and final time $T$. White circles indicate the observation locations.
@@
Figure 8: Time evolution of the relative RMSE for Example 2 under latent dimensions $n=16,20,24$. Solid curves indicate the mean over 10 independent runs, and the shaded regions represent the corresponding 95% confidence intervals.
@@
Finally, Tab. 2 summarizes the global relative RMSE $\mathrm{E}_{\mathrm{Rel},1:T}$ under different observation noise levels, together with the online computation time of the EnKF assimilation step. The results show that LAE-EnKF consistently achieves the best accuracy–efficiency trade-off. Compared with the standard EnKF, the proposed method substantially reduces computational cost while improving reconstruction accuracy. These results confirm that learning a stable linear latent representation and performing EnKF in the latent space can significantly enhance data assimilation performance for nonlinear PDE systems.
@@
Figure 9: Reconstructed fields and corresponding pointwise absolute error at the final time $T$ by LAE-EnKF for different latent dimensions $n=16$, $n=20$, and $n=24$.
@@
Figure 10: Comparison of reconstructed fields and corresponding pointwise absolute error at the final time $T$ for latent dimension $n=20$.
Table 2: Relative error $\mathrm{E}_{\text{Rel},1:T}$ and online computation time for Example 2.
@@
$\sigma=0.10$ 0.6885 0.3468 0.4130 0.1091
$\sigma=0.05$ 0.6705 0.2988 0.3024 0.0810
$\sigma=0.01$ 0.5119 0.2142 0.2213 0.0717
Time (s) 1174.10 3.77 0.05 0.04

### 29 Apr 2026 v2

@@ 5 Numerical Results
In this section, we present numerical experiments to assess the performance of the proposed LAE-EnKF in nonlinear data assimilation problems. The experiments are designed to evaluate three key aspects of the method: assimilation accuracy, robustness over time, and the benefit of enforcing linear dynamics in a learned latent space. All of the following numerical examples are carried out on a computing server with a NVIDIA GeForce RTX 3090 GPU. To enrich the information content of partial observations, we adopt a time-delay embedding strategy [41]. Specifically, a sequence of $L$ consecutive observations is concatenated to form an augmented observation vector,
@@
which is then processed by the observation encoder $\mathscr{E}_{\mathrm{obs}}$. This construction allows the latent observation embedding to exploit temporal correlations and improves identifiability by incorporating short-term dynamics.
@@
(14) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{E}_{1:K}=\sqrt{\frac{1}{KD}\sum_{k=1}^{K}\sum_{i=1}^{D}\big(\widehat{x}_{k,i}-x_{k,i}\big)^{2}},\quad\mathrm{E}_{\mathrm{Rel},1:K}=\frac{\mathrm{E}_{1:K}}{\sqrt{\frac{1}{K}\sum_{k=1}^{K}\left(\frac{1}{D}\sum_{i=1}^{D}x_{k,i}^{2}\right)}}.$}
@@
    EnKF: the standard ensemble Kalman filter, where both forecast and analysis steps are carried out directly in the physical state space (see Section 2.2).
@@
    AE-EnKF: an autoencoder-assisted EnKF where the encoder-decoder pair is used to reconstruct the state during the forecast step, while the EnKF update is performed in the physical space (see Algorithm 2).
@@
    DAE-EnKF: a latent-space EnKF method in which the latent dynamics are modeled by a general nonlinear mapping, and both forecast and analysis steps are carried out in the latent space (see Algorithm 3). This setting is similar in spirit to [29]. To ensure a fair comparison, the encoder and decoder architectures are identical to those used in the proposed LAE framework.
All methods use the same ensemble size, data normalization, and training configuration to ensure comparability. In particular, all state variables and observations are normalized to zero mean and unit variance prior to training and assimilation. Unless otherwise specified, we use the identity latent observation operator $\mathbf{H}=\mathbf{I}$ throughout. In all experiments, the encoder and decoder are implemented using convolutional neural networks (CNN), which are well suited for high-dimensional systems with spatial structure. The state encoder consists of three convolutional layers. Each layer uses a $4\times 4$ kernel with stride $2$, progressively reducing the spatial resolution while extracting multiscale features. A fully connected layer with LeakyReLU activations (negative slope $\alpha=0.1$) then maps the resulting features to the latent variables. The decoder mirrors this architecture and reconstructs the physical state using transposed convolutional layers. The latent linear transition operator $\mathbf{A}$ is learned as a single fully connected neural network without any activation function. This architectural choice explicitly enforces linear latent dynamics and is essential for maintaining structural compatibility with Kalman filtering. The observation encoder is implemented as a lightweight network composed of two one-dimensional convolutional layers with kernel size $3$, followed by a fully connected layer that projects the observations into the latent space. This design allows observations with different spatial layouts to be embedded consistently with the latent state representation. The loss weights in (3.2) are selected so that the reconstruction, prediction, latent consistency, and regularization terms contribute at comparable magnitudes during training. All networks are implemented in PyTorch and trained using the Adam optimizer with an initial learning rate of $10^{-3}$. Early stopping based on the validation RMSE is applied to mitigate overfitting.
@@ 5.1 Toy example
The system is governed by a single angular variable $\theta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}$, while the physical state is defined as
(15) $\mathbf{x}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}=W\begin{bmatrix}\cos(\theta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}})\\ \sin(\theta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}})\end{bmatrix}\in\mathbb{R}^{100},\qquad W\in\mathbb{R}^{100\times 2},\quad W_{ij}\sim\mathcal{N}(0,1),$
which embeds a two-dimensional rotational motion into a 100-dimensional ambient space. Here, $k=0,1,\dots,K$ denotes discrete time steps. The angular dynamics evolve according to
(16) $\theta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}+1}=\theta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}+\delta+\alpha\sin(2\theta_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}})+c\,\epsilon_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}},$
where $\delta=\pi/50$ is a constant rotation increment, $\alpha=0.01$ introduces weak nonlinearity, $c=0.01$ scales the process noise, and $\epsilon_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}\sim\mathcal{N}(0,1)$.
@@
[…]
where $\mathcal{H}:\mathbb{R}^{100}\to\mathbb{R}^{2}$ is a linear observation operator that randomly selects two state components. We simulate $N_{\mathrm{traj}}=500$ independent trajectories of length ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}=100$, with initial angles $\theta_{0}$ sampled uniformly from $\pi,\pi]$. The data are split into $90\%$ for training and $10\%$ for testing.
@@
Figure 4: Long-term prediction relative error for latent dimensions $n=2,3,4$. Predictions are obtained by recursively propagating the learned latent dynamics over time steps and decoding back to the physical space.
Finally, we evaluate data assimilation performance using LAE-EnKF and compare it with EnKF, AE-EnKF, and DAE-EnKF. All methods exhibit comparable computational costs within the same order of magnitude, run-time comparisons are therefore not reported. Tab. 1 reports the global relative RMSE $\mathrm{E}_{\mathrm{Rel},1:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}$ for latent dimensions $n=2,3,4$. Fig. 5 shows the evolution of the reconstruction error over discrete time steps for $n=2$, together with $95\%$ confidence intervals over ten independent runs. All methods outperform the standard EnKF. AE-EnKF attains low error in the initial stage, followed by oscillations and eventual error growth. This is due to its reliance on a one-step learned mapping, which lacks temporal consistency under repeated application, leading to error accumulation. In contrast, DAE-EnKF shows noticeable fluctuations, while LAE-EnKF produces a smoother and more stable error trajectory.
Table 1: Relative error $\mathrm{E}_{\text{Rel},1:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}$ for different methods with various values of $n$.
@@
[…]
Figure 5: Evolution of the relative RMSE over discrete time steps for different methods with latent dimension $n=2$. Solid curves show the mean over 10 independent runs, and shaded regions indicate the corresponding 95% confidence intervals.
[…]
@@ 5.2 Advection–diffusion–reaction equation
We consider a nonlinear advection diffusion reaction equation on the two-dimensional spatial domain ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}}=(x_{1},x_{2})\in[0,1]\times[0,1]$,
[…]
$u({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}},0)$ $=u_{0}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}}),$
subject to periodic boundary conditions. The diffusion coefficient and reaction parameter are set to $\mu=10^{-5}$ and $\alpha=0.8$, respectively. The advection velocity field is given by
$\mathbf{v}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}})=\bigl(\sin(2\pi x_{2}),\,\sin(2\pi x_{1})\bigr).$
@@
[…]
@@
$K_{\ell}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}}^{\prime})=\exp\!\left(-\frac{\sin^{2}(\pi(x_{1}-x_{1}^{\prime}))+\sin^{2}(\pi(x_{2}-x_{2}^{\prime}))}{2\ell^{2}}\right).$
Data assimilation is performed using sparse and noisy observations of the true solution. Measurements are collected at $5\times 5$ uniformly distributed sensor locations, resulting in an observation vector $\mathbf{y}_{k}\in\mathbb{R}^{25}$, corrupted by additive Gaussian noise with standard deviation $\sigma=0.01$. Observations are available every $\Delta t=0.05$ up to the final time $T=3$. Fig. 7 shows the true solution at the initial and final times, with sensor locations marked by white circles. The initial ensemble is generated by sampling spatial white noise followed by convolution with a two-dimensional Gaussian filter with standard deviation $3$. This procedure introduces spatial correlation and provides a common prior ensemble for all filtering methods.
@@
Figure 7: True solution fields of the advection–diffusion–reaction equation at the initial time (Left) and final time $T=3$ (Right). White circles indicate the observation locations.
To better understand the role of different components in the proposed framework, we first conduct an ablation study on the loss function (3.2). Here, we fix the latent dimension to $n=20$ and remove either the latent consistency term or the prediction term in (3.2). Fig. 8 shows the multi-step state prediction relative error $\frac{\|\mathscr{D}\mathbf{A}^{k}\mathscr{E}(\mathbf{x}_{0})-\mathbf{x}_{k}\|_{2}}{\|\mathbf{x}_{k}\|_{2}}$ and the latent consistency relative error $\frac{\|\mathbf{A}^{k}\mathscr{E}(\mathbf{x}_{0})-\mathscr{E}(\mathbf{x}_{k})\|_{2}}{\|\mathscr{E}(\mathbf{x}_{k})\|_{2}}$. The results indicate removing the latent consistency term significantly degrades the linear structure in the latent space, while removing the prediction term leads to reduced accuracy in the physical state. This confirms that the two terms play complementary roles in learning accurate and structured latent dynamics.
@@
Figure 8: Ablation study of the loss function (3.2) with latent dimension $n=20$. Left: relative state prediction error $\frac{\|\mathscr{D}\mathbf{A}^{k}\mathscr{E}(\mathbf{x}_{0})-\mathbf{x}_{k}\|_{2}}{\|\mathbf{x}_{k}\|_{2}}$. Right: relative latent consistency error $\frac{\|\mathbf{A}^{k}\mathscr{E}(\mathbf{x}_{0})-\mathscr{E}(\mathbf{x}_{k})\|_{2}}{\|\mathscr{E}(\mathbf{x}_{k})\|_{2}}$.
@@
[…]
Figure 9: Time evolution of the relative RMSE for Example 5.2 under latent dimensions $n=16,20,24$. Solid curves indicate the mean over 10 independent runs, and the shaded regions represent the corresponding 95% confidence intervals.
@@
Figure 10: Reconstructed fields (Top) and corresponding pointwise absolute errors (Bottom) at the final time $T=3$ by LAE-EnKF for different latent dimensions $n=16$, $n=20$, and $n=24$.
@@
Figure 11: Comparison of reconstructed fields (Top) and corresponding pointwise absolute errors (Bottom) at the final time $T=3$ for latent dimension $n=20$.
We next examine the effect of different latent observation operators. In addition to the choice $\mathbf{H}=\mathbf{I}$, we also consider both randomly generated observation operators and learned observation mappings in the latent space. For a fair comparison, the latent dynamical model obtained in the first training stage is kept fixed across all settings. Fig. 12 shows the time evolution of the relative RMSE under these different observation models, with the global error $\mathrm{E}_{\text{Rel},1:K}$ reported in the title. The results demonstrate that comparable performance is achieved across all observation operators, indicating that the proposed method is not sensitive to the specific choice of the observation model.
[…]
Figure 12: Time evolution of the relative RMSE for Example 5.2 under different latent observation operators: identity operator (Left), randomly operator (Middle), and learned operator (Right). Solid curves denote the mean over 10 independent runs, and shaded regions indicate the corresponding 95% confidence intervals. The global relative error $\mathrm{E}_{\text{Rel},1:K}$ is reported in each subplot.
Finally, Tab. 2 summarizes the global relative RMSE $\mathrm{E}_{\mathrm{Rel},1:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}$ under different observation noise levels, together with the online computation time of the EnKF assimilation step. The results show that LAE-EnKF consistently achieves the best accuracy–efficiency trade-off. Compared with the standard EnKF, the proposed method substantially reduces computational cost while improving reconstruction accuracy. These results confirm that learning a stable linear latent representation and performing EnKF in the latent space can significantly enhance data assimilation performance for nonlinear PDE systems.
Table 2: Relative error $\mathrm{E}_{\text{Rel},1:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}$ and online computation time for Example 2.
@@
$\sigma=0.10$ 0.7987 0.3371 0.4369 0.1058
$\sigma=0.05$ 0.7553 0.2855 0.3400 0.0780
$\sigma=0.01$ 0.7516 0.2344 0.2860 0.0633
Time (s) 2837.29 6.75 0.10 0.05